• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Solvignon

    Log entry time 13:20:40 on July14,2003

    Entry number 103830

    keyword=EPR is working fine.

    Jaideep and I decided to use the left side of the PI feedback box due to loose BNC connection.

    After adjusting the gain on the left box, we got a pretty good EPR spectrum.

    FIGURE 1